OpenStudy (anonymous):

Will medal!! Consider the following pair of equations: y = x + 4 y = –2x – 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (anonymous):

i need help with this problem

OpenStudy (kidrah69):

x+4=-2x-2 solve

OpenStudy (michele_laino):

hint: we can write: \[\Large x + 4 = - 2x - 2\]

OpenStudy (kidrah69):

all that was done was I plugged equation 1 into equation 2

OpenStudy (anonymous):

how do i do it with substitution?

OpenStudy (kidrah69):

that is substitution

OpenStudy (michele_laino):

more precisely: \[\Large x + 4 = y = - 2x - 2\]

OpenStudy (michele_laino):

you have to replace y with x+4, into the second equation